2p-3q=7 3p+2q=17, 3x+y=-1 x+y=7, 5e+d=12 3e-2d=2.
Answer by Boreal(15235)

2p-3q=7
3p+2q=17, multiply top by -3 and bottom by 2
-6p+9q=-21
6p+4q=34
13q=13
q=1
p=5
==
3x+y=-1
x+y=7
-x-y=-7
2x=08
x=-4
y=11
==
5e+d=12
3e-2d=2
10e+2d=24
13e=26
e=2
d=2